OpenStudy (anonymous):

Show that the function sin(x^3+7x+1) is continuous everywhere.

OpenStudy (anonymous):

The theorem the book tells us to use is, If the function g is continuous everywhere and the function f is continuous everywhere, then the composition f of g is continuous everywhere.

OpenStudy (zzr0ck3r):

so you need to show that sin(x) is continuous and x^3_7x+1 is continuous...do you need to show this with epsilon proof? do you have other theorems? like sin(x) is differentiable and thus continuous....

OpenStudy (anonymous):

Um, the theorem that the book tells us to use to show that the function is continuous is: If the function g is continuous everywhere and the function f is continuous everywhere, then the composition f of g is continuous everywhere.

OpenStudy (zzr0ck3r):

i get that, and we know that sin(x) and x^3-7x+1 are both continuous so the result follows having said that do we need to prove that sin(x) is continuous? Is this for a epsilon delta type calculus class?

OpenStudy (anonymous):

no, its just to show that the function is continuous.

OpenStudy (anonymous):

@zzr0ck3r r u still available?

OpenStudy (zzr0ck3r):

f(x) = sinx is continuous g(x) = x^3-7x+1 is continuous so by the theorem you stated f(g(x)) = sin(g(x)) = sin(x^3-3x+1) is continuous